### v1

# Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap

Abstract
In high-stakes, time-critical scenarios—such as emergency evacuation, first responder prioritization, and crisis management—decision-makers must rapidly choose among spatial targets, such as exits, individuals to assist, or areas to secure. Advances in indoor sensing and artificial intelligence (AI) can support these decisions by visualizing real-time situational data and AI suggestions on 2D maps. However, mentally mapping this information onto real-world spaces imposes significant cognitive load. This load can impair users’ ability to appropriately judge AI suggestions, leading to inappropriate reliance (e.g., accepting wrong AI suggestions or rejecting correct ones). Embedded visualizations in Augmented Reality (AR), by directly overlaying information onto physical environments, may reduce this load and foster more deliberate, appropriate reliance on AI. But is this true? In this work, we conducted an empirical study (N = 32) comparing AR see-through (embedded visualization) and 2D minimap in time-critical, AI-assisted spatial target selection tasks. Contrary to our expectations, users exhibited greater inappropriate reliance on AI in the AR condition. Our analysis further reveals that this is primarily due to over-reliance, with factors specific to embedded visualizations, such as perceptual challenges, visual proximity illusions, and highly realistic visual representations. Nonetheless, embedded visualizations demonstrated notable benefits in spatial reasoning, such as spatial mapping and egocentric spatial imagery. We conclude by discussing the empirical insights, deriving design implications, and outlining important directions for future research on human-AI decision collaboration in AR.
keywords:
Augmented Reality; Decision Making; Human-AI Collaboration; Situated / Embedded Visualizations
[…]
A participant wearing Apple Vision Pro performs spatial decision making tasks using two different visualizations: AR See-Through (Left) and 2D Minimap (Right). Both visualizations present AI-suggested targets, but differ in representations. The AR See-Through approach embeds information directly into the participant’s field of view, while the 2D Minimap offers a top-down abstraction.
Introduction
In high-stakes, time-critical scenarios – such as emergency evacuation [89, 50], first-responder prioritization [96], and crisis management [40] – decision-makers must rapidly select among multiple spatial targets (e.g., exits, individuals requiring assistance, or areas to secure). We refer to these decision-making tasks, which require integrating information about spatial environments with external data, as Spatial Decision-Making. Making the wrong choice in these tasks can waste valuable resources or endanger lives [31]. Recent advancements in indoor sensing and artificial intelligence (AI) offer new avenues for supporting such decisions by visualizing real-time situational data and AI suggestions in spatial context [59, 44].
However, the way this information is presented can significantly impact user behavior and decision quality [24]. Existing methods [45, 92] commonly employ 2D maps to display data and AI suggestions, requiring users to mentally map and connect these digital cues onto physical spaces. According to spatial cognition research [38], this process – known as reference-frame translation – can impose significant cognitive load. Under time pressure, users may struggle to critically assess AI outputs, resulting in inappropriate reliance [74]: either trusting flawed AI suggestions or rejecting correct ones. This concern is especially pressing in emergency contexts, where every second counts.
Augmented Reality (AR) promises a more intuitive approach by directly visualizing situational data and AI suggestions onto the physical environment. These visualizations – often referred to as embedded visualizations [87, 42, 8, 69] – can reduce cognitive load by eliminating the need for constant reference-frame translation. Intuitively, embedded visualizations might foster more deliberate and appropriate reliance on AI. Yet whether embedded visualizations genuinely mitigates inappropriate reliance on AI in spatial decision-making remains unclear, and little existing work has examined how embedded visualizations in AR affects human-AI reliance in spatial tasks. As AR and AI continue to mature and increasingly integrate [32], this work aims to provide researchers and practitioners with initial evidence on how embedded visualizations impact human-AI reliance in spatial decision-making.
To this end, we present an empirical study (N=32) comparing an AR see-through visualization (Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap Left) against a 2D minimap (Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap Right) in a time-critical, AI-assisted spatial decision task. Following previous work on AR visual systems for indoor navigation [59, 91, 90], we selected the AR see-through visualization because it naturally embeds information within a large-scale indoor environment. In the tasks, participants needed to select one of four coffee machines distributed across a two-floor building based on multiple criteria – walking distance and queue length – to obtain coffee as quickly as possible. An AI with 75% accuracy was provided to suggest the optimal target, and both visualizations were experienced using an Apple Vision Pro (AVP), a state-of-the-art AR headset.
Contrary to initial expectations, participants exhibited greater inappropriate reliance on AI with the AR see-through visualization, suggesting that new perceptual and cognitive challenges arise when digital cues are spatially embedded in large or complex physical environments. Further analysis of quantitative data (accuracy, response times, reliance metrics, and self-reported confidence) reveals that the inappropriate reliance primarily manifested as over-reliance – blindly accepting AI suggestions. Through qualitative interviews, we identify several key factors that contributed to the over-reliance, such as perceptual challenges in AR, visual proximity illusions, and heightened trust in “embodied” AI suggestions. On the other hand, our findings also highlight embedded visualizations’ strength in multiple spatial reasoning tasks, including improved spatial mapping (as evidenced by quantitative data) and support for egocentric spatial imagery (as reported by participants). We conclude by outlining the lessons learned, design implications, and promising future directions for better harnessing these AR strengths to foster human–AI collaboration in spatial decision-making.
1 Related Work
Human-AI Decision Making. Human-AI decision making refers to collaborative decision processes where a human decision-maker interacts with an AI system to make a choice [3] . A common goal in these systems is to achieve complementary performance, where the human-AI team performs better than either the human or the AI alone [3] . However, empirical studies have repeatedly found that such complementary performance is rarely observed (e.g., [5, 11]) . One widely reported barrier is inappropriate reliance — the inability of human users to correctly accept AI suggestions when they are right, or reject them when they are wrong [3] .
Unlike trust in AI, which reflects users’ subjective feedback, reliance objectively measures whether a participant accepts or rejects the AI’s suggestion [3, 49] , and has therefore been widely studied. To foster appropriate reliance, researchers have proposed providing explanations of the AI’s suggestions [67]. Unfortunately, several studies found that explanations have often failed to deliver their intended benefits, but actually increase over-reliance – a phenomenon where users uncritically accept AI suggestions, even when they are suboptimal [3]. Drawn on the dual-process model of cognition [12, 22], initial efforts attributed this over-reliance to a lack of cognitive engagement with the explanation, because people often rely on heuristics (Type 1) to judge the AI’s suggestions, rather than engaging in deeper analysis (Type 2). Buçinca et al. [10] introduced cognitive forcing functions that encourage users to think more critically about the AI’s suggestions. Vasconcelos et al. [79] proposed a cost-benefit framework that formalizes when users choose to engage with AI explanations. Recently, Guo et al. [28] developed a decision-theoretic framework to measure AI reliance by separating reliance behavior from cognitive limitations in signal interpretation.
In the context of visualization research, automated methods to assist humans in making decisions with data have long been practiced. While a large body of work has focused on developing visual analytics systems [65], relatively fewer studies have investigated how visualizations affect human-AI collaboration. Among these studies, Yang et al. [93] explored how visual explanations can foster appropriate trust in AI systems; Morrison et al. [51] investigated how different human explanation strategies might impact users’ reliance on AI in visual decision-making; Gaba et al.[24], Wang et al.[83] , and Wall et al. [81] each focused on how interface design influences users’ perceptions of model bias and performance, highlighting the powerful role of visual presentation in shaping trust on AI systems. Recently, Ha et al. [29] found that in AI-guided visual analytics systems, users were more inclined to accept suggestions when completing a more difficult task despite the AI’s lower suggestion accuracy. Zhao et al. [97] and Reyes et al. [63] both investigated how visualizing uncertainty in model outputs influences reliance and trust. Taken together, these studies underscore that visual presentation profoundly influences how users evaluate AI suggestions, rely on them, and ultimately make decisions.
Unlike all these existing works, which evaluate desktop-based human-AI collaboration, our study contributes an initial empirical investigation of the impact of AR visualizations on human-AI collaboration in spatial decision-making tasks. The differences extend beyond display technique (2D monitor vs. AR) to the task environment (desktop vs. physical environment) and to cognitive processes (e.g., general vs. spatial cognition), opening new questions that merit dedicated study.
Visualization-Aided Decision Making. Within the field of visualization, research on decision-making typically falls into two major categories: understanding how humans make decisions when using visualizations (e.g., [54, 15, 4]), and designing interactive systems that support decision-making (e.g., [53]).
The former body of work has drawn from theories in cognitive psychology to model and explain the mechanisms by which visualizations influence decision-making. For example, Padilla et al. [54] proposed a cognitive framework grounded in dual-process theories, illustrating how visualizations engage both intuitive (Type 1) and analytical (Type 2) reasoning processes in different contexts. Bancilhon et al. [2] expanded on this perspective by advocating for the integration of cognitive models in evaluating visualization effectiveness. A parallel line of work investigates how visualizations interact with human cognitive biases. Dimara et al. [15] introduced a task-based taxonomy of cognitive biases relevant to visualization, while Wall et al. [82] explored how interaction traces might mitigate such biases. Bearfield et al. [4] further demonstrated that the same data visualization can elicit diverging interpretations. Research on uncertainty representations also plays an important role in understanding decision-making: Kale et al. [36] showed that design choices, such as emphasizing means, can bias effect size judgments, while Fernandes et al. [23] found that quantile dotplots can significantly improve real-world decisions in uncertain scenarios.
The latter body of work focuses on developing interactive visualization tools to support decision-making. Some of these systems are domain-agnostic and designed to facilitate specific decision processes, such as inspecting rankings [80] and comparing alternatives based on multiple, often conflicting attributes [56] in Multi-Criteria Decision-Making. Another line of work targets domain-specific decision contexts, such as medical treatment planning [39] , urban policy [84] , and more [13] . These systems typically incorporate domain-specific constraints to guide and contextualize the decision process. A recent survey by Oral et al. [53] offers a comprehensive analysis of decision-focused visualization tools and highlights the opportunities and challenges in supporting all stages of the decision-making process.
While decision-making is widely recognized as a core goal of visualization, Dimara and Stasko [16] argue that it remains underrepresented in visualization research and task taxonomies. Brumar et al.[9] recently proposed a typology to better describe decision-making tasks in visualization. Despite this growing interest, decision support for spatial tasks remains an underexplored yet promising area [26]. Our study takes a first step toward investigating spatial decision support, aiming to provide a humble baseline and reference point for future research.
Spatial Cognition and Spatial Decision-Making. Spatial cognition encompasses the mental processes involved in perceiving, encoding, and reasoning about spatial environments [77]. A fundamental aspect of spatial cognition is the use of reference frames—mental models that help individuals locate and orient themselves in space [77, 76] . Two commonly described reference frames are egocentric, where objects are located relative to the observer’s current position and orientation (i.e., “to my left or right”), and allocentric, where the environment is represented independently of the observer’s position (e.g., map-like or object-to-object relationships) [38] . People switch flexibly between these frames, but doing so can introduce significant cognitive load [76] , particularly when tasks require precise mental transformations, such as rotating or scaling spatial representations in mind.
Spatial decision-making requires integrating information about physical environments with external data. To make such decisions with AI support, decision-makers must reconcile digital information with real-world contexts. With traditional 2D map-based visualizations, this process often involves switching between reference frames, which is cognitively demanding and can affect interpreting or verifying AI suggestions. By contrast, AR can display information directly in the physical environment, thereby eliminating the spatial mapping overhead [6]. Yet, to the best of our knowledge, few studies have investigated the effect of AR visualizations on human-AI decision making. Moreover, how to best design effective AR visualizations for spatial decision-making remains an open question. These gaps motivate our empirical investigation into how AR visualizations might influence human-AI collaboration in spatial decision-making.
Situated, Embedded, and AR Visualizations. Different from Virtual Reality (VR), AR can directly visualize information within physical environments, benefiting a broad range of applications that require data-driven or computational support in situ [86, 8]. The visualization community has extensively explored various aspects of AR visualization, including design patterns [69, 42], techniques [98], empirical evaluations [58], and applications [46]. More broadly, the concept of visualizing information in its physical context is known as Situated Visualization [69, 42], for which AR is one realization. Depending on the spatial relationship between the visualization and its physical referent, Willett et al. [87] distinguish between situated (co-located in physical space) and embedded (overlaid or integrated with the referent). Based on this definition, our study compares two AR visualization approaches—a situated 2D minimap and an embedded AR see-through overlay—to investigate whether the embedded approach can foster greater human reliance on AI by eliminating the need to mentally map data from a separate panel onto the physical environment.
AR Visualization for Spatial Decision-Making. Our particular interest lies in leveraging AR to support indoor spatial decision-making under time pressure, a crucial area with high-stakes applications such as emergency evacuation [89, 68], first response operations [96], and security management [73]. Despite its importance, research specifically addressing time-constraint spatial decision-making in AR remains limited. Perhaps the closest related work involves AR-based evacuation assistance, which provides users with dynamic, context-sensitive navigation cues to safely guide them toward exits during emergencies [96]. For instance, Wächter et al. [89] investigated AR-guided evacuation in buildings, showing that adapting AR visualizations to environmental hazards can improve evacuation efficiency and safety. Similarly, Sharma et al. [68] developed AR modules to enhance situational awareness and cognitive mapping during evacuations in buildings.
Beyond emergency scenarios, a broadly related line of work explores AR-based indoor navigation, where users typically follow predefined routes rather than making urgent, real-time spatial decisions. Most AR navigation techniques fall into two categories: annotation-based methods (e.g., arrows) and see-through visualizations. Annotations are arguably the most common AR navigation technique, as seen in applications like Google Maps. Numerous studies (e.g., [43, 17]) have compared annotation-based AR navigation to traditional 2D minimaps across various contexts. On the other hand, AR see-through methods offer benefits like a global view but have received less attention, partly due to technical challenges requiring the development of digital twins of real-world environments. Representative work in this direction includes Xu et al. [91, 90] who conducted comparative studies of AR see-through visualizations for indoor wayfinding, demonstrating that egocentric perspectives significantly enhanced navigation efficiency, reduced cognitive load, and improved spatial awareness.
Drawing inspiration from these works, we adopt AR see-through visualizations to present spatial information within indoor environments. Unlike previous studies, however, our research specifically investigates how AR see-through techniques can support users in making spatial decisions with AI assistance – a topic that has not yet been explored in depth. Our study thus aims to provide an initial reference point for future research into AI-aided decision-making in AR contexts.
2 Study Rationale
This section defines the scope and abstraction of the spatial decision-making task, and summarizes the rationale behind our choices of visualizations, decision supports, and study hypotheses.
2.1 Task Motivation and Abstraction
Spatial decision-making has long been studied under the broader umbrella of wayfinding in spatial cognition (e.g., [21, 85]). Foundational work by Siegel and White [70] conceptualizes wayfinding as a decision process involving three forms of spatial knowledge: landmark, route, and survey. Motivated by real-world applications, our study particularly focuses on target selection (i.e., landmark). In high-stakes scenarios, such as emergency response, indoor search-and-rescue, and security surveillance, spatial decision-making often involves identifying or selecting a target before proceeding with an action (e.g., navigation or intervention). Although AR systems have frequently been investigated for supporting the action phase (e.g., route guidance [90]), their potential to enhance the preceding decision-making step remains underexplored.
To address this gap, we abstract real-world scenarios into a domain-agnostic spatial target selection task, defined by the following core attributes derived from the literature:
  • Space - Large-scale Indoor Environment: These tasks often take place in physical spaces larger than the immediate space around a person (e.g., buildings or open areas) [25, 27]. Tversky [76] characterize such spaces as space of navigation. In this study, we focus on indoor building environments, given their prevalence across numerous application domains.
  • Time - Sensitive: Many high-stakes spatial tasks impose stringent time constraints, requiring users to make optimal decision as soon as possible [52]. In this study, we focus on time pressure scenarios which particularly benefit from decision support systems.
  • Data - Static and Dynamic: Users typically can assess both static information (e.g., building layouts) and dynamic updates (e.g., occupant counts, hazard alerts) from sensors to evaluate each target [1].
  • Decision - Multi-Criteria: These tasks frequently require decisions among multiple targets or locations distributed across rooms, floors, or areas [40]. Success often involves balancing or prioritizing several factors, such as response time, travel distance, and urgency [57].
By focusing on these domain-agnostic factors, we can investigate how AR-based decision support influences spatial decision-making without limiting our findings to one specific use case.
2.2 Visualization Methods: Minimap vs. AR See-Through
Given that spatial decision-making inherently involves interpreting spatial relationships, an effective decision support tool must present relevant data within its spatial context. Thus, we select map-based visualizations as the foundational interface for our study. We are particularly interested in how different visualization paradigms – traditional 2D vs. embedded – influence users’ decision-making processes. Therefore, we choose two representative visualization methods: a conventional 2D minimap and an AR see-through visualization. Below, we discuss the rationale behind choosing these two methods and briefly highlight alternative visualization approaches.
2D Minimap. 2D minimap is perhaps the most widely adopted map-based visualizations on traditional flat-screens. It presents the spatial environment as a simplified, abstracted, top-down representation, upon which additional data can be annotated (e.g., Google Maps [^1] https://maps.google.com). For indoor environments spanning multiple floors, 2D minimap often employs separate layers for each floor, indicating spatial relationships through key landmarks such as stairways or elevators. 2D minimap is well-established and require minimal training, making them a suitable baseline for comparison. In our study, we implement a minimap (Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap Left) that shows the spatial distribution of multiple targets and their associated data (i.e., number of people in line) using glyphs.
AR See-Through. Given our task focuses on large-scale indoor environments, we select AR See-Through (Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap Right) as our embedded visualizations, which overlays a 1:1 scale 3D map on the real-world. It allows users see the interior layouts or objects behind walls, as if they had “X-ray” vision [47], and has been widely used in applications like navigation [91] and construction [72]. For our study, we follow previous similar research [59, 90, 91] that visualize a building’s indoor layout, multiple targets (represented as cubes), and their associated data (i.e., the number of people in line) using virtual avatars.
Other alternatives. We have also considered other map-based alternatives. For example, the 3D minimap provides a three-dimensional representation of the environment and can be implemented on either flat-screen or AR displays. However, 3D minimaps typically rely heavily on interactions such as panning, tilting, zooming, or rotation to fully leverage their advantages. Because our current investigation prioritizes evaluating the impact of visual representations themselves (without the influence of interactive complexity), we intentionally avoid these interactions in our study. Thus, we reserve exploration of interactive 3D minimaps and other advanced visualization methods for future research.
2.3 Visual Decision Supports
A recent survey by Oral et al. [53] highlights a broad spectrum of visual decision support tools, ranging from basic information visualizations to advanced AI-based suggestions [41]. For this study, we specifically choose to implement two fundamental yet representative forms of decision support across both the 2D minimap and AR see-through:
  • Target Visualizations that presents candidate targets and associated real-time data (e.g., occupancy counts, hazard alerts) on the map, without path suggestions or explanatory text.
  • AI-Suggested Optimal Target that highlights a single optimal target based on predefined criteria, with no accompanying explanation. Similar to all AI system, suggestions may not always be accurate. We detail our adjustment of its performance based on prior works [74] and pilot studies in Sec. 3.7.
Following practices established in prior research [79], we deliberately constrain our decision support to these fundamental features for several reasons: First, target and real-time data visualization represent core functionalities broadly used in existing decision-support tools. Second, simple AI suggestions enable direct investigation of user reliance on AI assistance under different visualizations, without the additional cognitive complexity and potential confounding effects introduced by detailed explanations [35].
While more advanced support – such as comparative visualizations, AI-generated explanations, or uncertainty displays – may offer added benefits, we intentionally focus on these fundamental forms to isolate the role of embedded visualizations in spatial decision-making and AI reliance. This controlled setup provides a baseline for future studies exploring richer decision-support designs.
2.4 Hypothesis
According to dual-process theory, human decision-making can be categorized into two distinct cognitive systems: rapid, heuristic-driven Type 1 thinking, and slower, analytical Type 2 thinking [12]. A widely accepted distinction between these systems is that Type 2 thinking demands significantly more cognitive resources than Type 1 [12].
Under conditions of high time pressure, users often lack sufficient cognitive bandwidth to analytically evaluate AI-generated suggestions, potentially resulting in inappropriate reliance – either rejecting correct suggestions or accepting suboptimal ones [48]. This risk is especially pronounced with 2D minimap visualizations, as these interfaces require users to mentally map the information to their corresponding physical environment, imposing significant cognitive load due to reference-frame switching [30]. In contrast, embedded visualizations can significantly reduce or eliminate these spatial mapping demands by directly fusing information within the physical environment, thus potentially fostering more appropriate reliance on AI suggestions.
Based on this rationale, we hypothesize that in AI-aided spatial decision-making, compared to the 2D minimap:
  • H1
    AR see-through promotes more appropriate reliance on AI.
  • H2
    AR see-through improves overall task performance.
  • H3
    Response times is comparable across user interfaces.
3 Study Design
We conducted a 2 $\times$ 2 within-subjects study to investigate whether AR see-through can foster more appropriate reliance on AI in spatial decision-making. The two independent variables were the type of visualizations – AR (AR see-through) vs. Minimap (minimap) – and the availability of AI assistance (NoAI vs. AI). The study design is detailed below, with fixed values based on a 13-participant pilot study.
3.1 Task Description
We designed a realistic indoor building scenario, where participants acted as busy employees. The goal was to simulate a time-critical spatial decision-making scenario that closely resembles everyday scenarios to the participants. In each trial, participants were tasked with selecting one coffee machine from four available options within a two-floor building within 20 seconds. The objective was to choose the machine that would allow them to obtain a cup of coffee in the shortest possible time. Optimal decision-making in this task depended on two key factors:
  • Walking Distance — the spatial location of each coffee machine relative to the participant’s current position.
  • Queue length — the number of people currently waiting in line at each machine.
Participants accessed information about both the machine locations and queue lengths via one of two visualizations, and in some trials, they also received AI suggestions.
3.2 Spatial Arrangement
The task was set in a real two-floor indoor building spanning roughly 80m $\times$ 80m. In each trial, the participant and the four coffee machine targets were positioned based on a predefined spatial arrangement.
Target and Participant Locations. To simulate diverse spatial scenarios and mitigate learning effects, we varied the distribution of the four coffee machines across trials based on two spatial factors identified during the pilot testing:
  • Floor Relationship – same or cross: Targets were either located on the same floor as the participant or on a different floor (cross). This factor has a strong impact on Minimap conditions as cross-floor targets made estimating walking distances more challenging.
  • Proximity – close or far: Targets were categorized as either close or far in direct proximity (instead of walking distance) to the participant’s location. This factor has an impact on AR conditions, as distant targets were harder to perceive.
The combination of these two factors resulted in four difficulty levels, ranging from Same+Close (easiest) to Cross+Far (most difficult). To maintain perceptual consistency, all four targets within a trial shared the same floor and proximity level; therefore, no one target is perceptually significantly different from others. The participant’s location was then positioned at the center of a circle passing through all four targets.  Figure 2 illustrates an example of the four difficulty levels.
Grouping Arrangements by Participant Location. Ideally, both participant and target locations would vary each trial. However, moving participants between locations would introduce significant overhead and fatigue. To address this, we grouped arrangements by participant location. Specifically, we selected eight participant locations evenly distributed in the building (Figure 1). For each location, we designed four spatial arrangements, each representing one difficulty level, resulting in $4\times 8=32$ unique spatial arrangements. As a result, the arrangements were naturally grouped into sets of four, each set sharing the same participant location but differing in target distributions. Example of one group is shown in  Figure 2.
Target Optimality. Target optimality was determined by two metrics: walking distance from the participant and its queue lengths (i.e., number of people waiting). For simplification, we estimate the total time to obtain coffee from a target using this formula:
$\text{Time}_{\text{coffee}}=\frac{\text{Distance}_{\text{walk}}}{\text{Speed}_{\text{walk}}}+\text{Number}_{\text{people}}\times\text{Time}_{\text{wait}}$
Based on the pilot studies, we empirically set the walking speed at 1m/s and waiting time per person in line at 15s. These parameters were explicitly communicated to participants in the study.
Since the spatial arrangement of targets is fixed, we manipulated the queue lengths to clearly differentiate target optimality. Based on prior research [74], we intentionally included suboptimal choices to effectively evaluate potential participant over-reliance on AI suggestions. Each trial thus included one optimal, two suboptimal, and one worst choice, each differing by at least 10-second $\text{Time}_{\text{coffee}}$. Detailed calculations are provided in the supplementary materials.
3.3 Conditions
With a within-subjects design, each participant completing trials under four conditions: Minimap+NoAI, Minimap+AI, AR+NoAI, and AR+AI. The order of conditions was counterbalanced across participants using a Latin square design to minimize order effects and learning biases. Our primary focus was on participants’ behavior in the AI conditions, while the NoAI conditions served as baselines to better understand user decision-making patterns.
Implementation and Apparatus. Following similar study [10], a simulated AI was used to provide decision suggestions in applicable trials. Details of the simulated AI setup are provided in Sec. 3.7. We developed the minimap based on high-precision CAD models provided by the building manager. The AR see-through visualization was developed using ARKit  [^2] https://developer.apple.com/augmented-reality/ to scan the building and construct a 3D digital twin in Unity  [^3] https://unity.com/, with ARAnchor Manager ensuring proper registration of the digital model with the physical environment. The system is available at our public repository http://public-when-ready.
All trials were completed using an AVP  [^4] https://www.apple.com/apple-vision-pro/ to eliminate device-related differences and provided a high-fidelity immersive experience. Participants were required to have normal vision, use contact lenses, or can use the prescription inserts provided by us.
Figure 1: The 8 participant locations in the study, distributed across two floors of the building. Participants were instructed to complete 4 trials at each location, following the order indicated by the numbered labels. The letter “L” is an abbreviation for location.
Figure 2: An example spatial arrangement for a single participant location, illustrating four distinct difficulty levels: same+close, same+far, cross-floor+close, and cross+far. The participant’s location is marked in red, and colored stars indicate target locations in each difficulty level.
3.4 Participants
We recruited 32 participants (16 male, 16 female; ages 18–28) from a university via email, word of mouth, and posters. The study lasted approximately one hour, and participants received a \$15 Amazon gift card as base compensation. To incentivize accurate performance, participants earned an additional \$0.20 for each trial in which they selected the optimal target, with the bonus capped at \$20.
To minimize bias related to familiarity with the experimental setting, we pre-screened participants to ensure limited exposure to the building used in the study. Specifically, 19 participants (9 male, 10 female) had never visited the building, 12 participants (6 male, 6 female) had been in the building between one and five times, and 1 participant (male) had visited the building between five and ten times.
3.5 Procedures and Design
We used the following full-factorial within-subject study design with Latin square-randomized order of the conditions to minimize learning and fatigue effects:
[…]
Participants were split evenly between the 4 Latin square-randomized condition orders. The orders of the spatial arrangements and participant locations were kept consistent across all participants to control for any confounding effects related to environmental layout (see Figure 3).
Figure 3: One example of Latin square-randomized orders across the four conditions. The order of visualizations (Minimap, AR) and AI (AI, NoAI) conditions would be counterbalanced. Each row represents two groups of spatial arrangements within the same condition, each covering four difficulty levels. Each rectangle represents a single trial. Labels (“opt”, “sub”, “worst”) indicate which target the AI suggests in that trial.
Procedure Overview. Prior to the experiment, participants were required to take an elevator ride from the ground floor to the experiment floor (2nd floor) and wait at the elevator door for the research team. This procedure was designed to minimize any prior exploration or accumulation of spatial knowledge about the experimental environment. The study lasted approximately one hour in total:
  • Introduction (10mins): Participants first provided informed consent and were explicitly informed about the performance-based incentives. Next, participants received a detailed tutorial that explained the study motivation, the AVP’s calibration and usage, and task procedures.
  • Training Tasks (10mins): Participants completed four training trials – one per condition – to familiarize themselves with the task and visualizations. In these trials, after participants made their selections, the optimal target was explicitly highlighted to illustrate optimal decision-making. Participants were encouraged to ask questions and were informed they could withdraw at any point.
  • Actual Tasks (30mins): Participants completed 32 trials, grouped into eight blocks (one block per standing location, see Figure 3). Within each block, they remained physically stationary and completed four trials (one per spatial difficulty). Each condition was tested across two consecutive participant locations (i.e., eight trials per condition). Breaks were allowed during the study.
  • Post-study Interview (10mins): Participants were interviewed with a list of open-ended questions (attached in supplements).
Detailed Trial Procedure. Each trial proceeded through the following steps: The experimenter escorted the participant to the designated starting point and helped them put on the headset. When ready, the experimenter used a controller to initiate the trial, activating the visualization and a 20-second countdown timer. Participants verbally announced their selection (e.g., “Target A”) as soon as they identified the optimal target. Then, the experimenter stopped the timer, automatically clearing the visualization. Participants rated their confidence in their choice using a 1 (Strongly unconfident) to 7 (Strongly confident) scale. Lastly, participants physically pointed toward the actual spatial location of their chosen target. This process was repeated for each trial.
Figure 4: Classification of reliance as appropriate, under-, and over-reliance based on the alignment between AI suggestions and user decisions.
3.6 Measures
We recorded each session through audio and first-person view video. For every trial, we documented and calculated the following measures based on prior research [74, 66]:
  • Accuracy: We defined accuracy as the average point score across trials in each condition, ranging from 0 to 1. Following previous research [74], participants earned points according to their choice quality: optimal choice (1 point), suboptimal choices (0.5 points), and the worst choice (0 points). Each trial contained exactly one optimal target, two suboptimal targets, and one worst target.
  • AI Reliance: Following previous studies [93], we categorized participant reliance on AI as either appropriate or inappropriate (Figure 4). We classified AI reliance as appropriate when participants either followed correct AI suggestions or overrode incorrect ones with better choices. Inappropriate reliance included: (1) over-reliance: accepting suboptimal or worst suggestions; and (2) under-reliance: rejecting correct suggestions in favor of a worse option.
  • Response Time: We measured the duration from the start of each trial to the moment the participant verbally announced their choice.
  • Point Error: This measure indicates whether participants correctly pointed to the actual physical location of their selected target. More pointing errors suggest that the visualization falls short in supporting spatial mapping from the digital to the physical space.
  • Confidence: Participants rated their confidence in their decision on a Likert scale from 1 (Strongly unconfident) to 7 (Strongly confident).
3.7 Using Simulated AI
Similar to prior works [74, 10], we used a simulated AI with an average accuracy of 0.75 to ensure controlled experimentation. In each AI condition, the AI suggested the optimal target in five trials, a suboptimal target in two trials, and the worst target in one trial (Figure 3). This setup yields an average accuracy of 0.75 = $\frac{5\times 1+2\times 0.5+1\times 0}{8}$. The order of these suggestions was randomized and counterbalanced across participants, ensuring that optimal, suboptimal, and worst suggestions were evenly distributed across all trials and participants.
Mitigating effects of trust in AI. Following previous work [3, 10], we took steps to neutralize trust effects so that observed differences in reliance could be attributed primarily to the visualizations. Specifically, we avoided anthropomorphizing the AI [33], consistently referring to it as “the AI”. We also described AI outputs as “suggestions” to counteract perfect automation schemas [19] and emphasize that these suggestions could be incorrect. Furthermore, we do not provide decision feedback for each trial [14], and any information about the AI’s accuracy [95].
4 Results
We report both the quantitative and qualitative results from the study.
4.1 Quantitative Results
[…]
Figure 5: Summary of quantitative results across conditions. Metrics include (A) accuracy, (B) task completion time, (C) completion time excluding initial search, and (D) confidence ratings. Comparisons are shown between AI-assisted and non-AI conditions for both Minimap and AR visualizations, with significance levels indicated.
4.1.1 Accuracy
[…]
Further analysis suggests that AR+AI under-performance stems from inappropriate reliance on AI. The AI system alone achieved a baseline accuracy of 0.75, while human participants in AR+NoAI achieved $Mdn=0.66$. Surprisingly, participants in the AR+AI condition ($Mdn=0.70$) still performed below the AI baseline, suggesting inappropriate reliance on AI. This aligns with findings in prior empirical studies [3]. In contrast, performance in both Minimap+NoAI ($Mdn=0.88$) and Minimap+AI ($Mdn=0.88$) exceeded the AI baseline, indicating that participants were able to appropriately rely on AI by identifying and overriding poor suggestions.
[…]
4.1.2 AI Reliance
Results from the accuracy analysis indicated that participants relied on AI inappropriately in the AR+AI condition. Quantitative analysis confirms this pattern: we observed a significant difference between Minimap+AI and AR+AI, with participants showing less appropriate reliance on AI in AR+AI (Figure 6a). This finding again contradicts our second hypothesis H2.
Inappropriate reliance on AI can be either over-reliance or under-reliance. To examine which pattern dominated participants’ interactions with AI in AR+AI, we separately analyzed trials of over-reliance and under-reliance. Figure 6b and c depict the average occurrences of over- and under-reliance, respectively, per eight trials.
To interpret these numbers, we compared observed behaviors to a random-selection baseline. With four possible options per trial, random selection would result in an expected average of 0.75 over-reliance instances and 4.25 under-reliance instances per eight trials (see supplementary material). Compared to this baseline, participants in the AR+AI condition displayed fewer instances of under-reliance ($M_{\text{{AR}}}=1.63<M_{\text{random}}=4.25$), but notably more instances of over-reliance ($M_{\text{{AR}}}=1.41>M_{\text{random}}=0.75$). In contrast, participants in the Minimap+AI condition exhibited fewer instances of both under-reliance ($M_{\text{{Minimap}}}=0.84<M_{\text{random}}=4.25$) and over-reliance ($M_{\text{{Minimap}}}=0.72<M_{\text{random}}=0.75$).
These findings indicate that inappropriate reliance in the AR+AI condition was primarily driven by over-reliance, consistent with our earlier accuracy analysis (Sec. 4.1.1).
Figure 6: Results of AI reliance across visualizations. Metrics include (A) appropriate reliance, (B) under-reliance, and (C) over-reliance, comparing Minimap+AI and AR+AI conditions. Random-selection baselines (in red) are indicated.
4.1.3 Response Time
We computed the average response time across conditions and observed significant effects (Figure 5b). Participants responded faster in Minimap+AI than AR+AI ($\Delta_{M}=\text{-0.94},p=\text{0.01}$) and in Minimap+NoAI than AR+NoAI ($\Delta_{M}=\text{-2.00},p<\text{0.001}$). These results contradict our null hypothesis H3.
Response Time Excluding Initial Search. We observed that participants in the AR conditions spent additional time to visually searching the spatially distributed targets—especially those outside their immediate field of view. In contrast, in the Minimap conditions, all targets were presented in a 2D overview, making search overhead negligible. Prior decision-making models (e.g., Simon’s model [71], the drift–diffusion model [61], the leaky competing accumulator model [78]) and empirical studies (e.g., [62, 64]) suggest that actual decision making begins only after sufficient information has been gathered. This motivates us to exclude the initial search overhead in the AR conditions to better compare its decision making phase with the Minimap conditions.
Figure 7: (a) and (b) show the direct and walking distance to the orange target; (a) and (b) also indicate the start and end moment of searching for the blue target.
Following prior visualization-decision work [55], we consider that meaningful decision making start only after at least two options have been identified. Thus, for each trial in the AR conditions, we excluded the initial search time required to discover the first two distinct targets. Specifically, we defined two search segments to exclude: S1-from the start of the trial to the first frame when the first target entered the user’s field of view; and S2-from the frame when the user turned away from the first target (e.g., Figure 7a) to the frame when the second target appeared (e.g., Figure 7b). If a target was already visible at trial start, or if the first and second targets appeared simultaneously, the corresponding segment (S1 or S2) was set to zero. This definition provides a strict lower bound on search time, as participants may still require additional head movement to center the target in view. Two co-authors independently annotated these segments frame by frame from the first-person recordings to ensure inter-rater reliability. Discrepancies were discussed and resolved through consensus.
After subtracting the annotated initial search time from the response time, we observed a different pattern (Figure 5c). Participants in AR+AI spent less time than those in Minimap+AI with a significant effect ($\Delta_{M}=\text{-0.85},p=\text{0.01}$), indicating that, with AI assistance, decisions were made more quickly with the embedded visualization. No significant difference was observed between Minimap+NoAI and AR+NoAI, suggesting that, without AI assistance, participants spent comparable time making decisions across visualizations In summary, although raw response times were longer in AR conditions, a detailed breakdown reveals that this extra time wasn’t spent on the decision making. This leads to interesting implications that we will further discuss in Sec. 5.4.
4.1.4 Post-Trial Pointing Error Rate
[…]
4.1.5 Self-Reported Confidence
We also examined participants’ self-reported confidence ratings across the four conditions. This measure reflects how confident participants were that they had selected the optimal target in each trial. Our analysis revealed no significant differences across conditions (Figure 5d), suggesting that participants felt equally confident in their decisions regardless of visualization type or AI assistance.
4.2 Qualitative Results
We conducted semi-structured post-study interviews to capture participants’ strategies, rationales, and motivations behind their decision-making across all conditions. Participants are referred to as P1–P32. We analyzed the interviews using reflexive thematic analysis [7], identifying themes both inductively (bottom-up) and deductively (top-down). Below, we summarize the key themes that emerged from the analysis.
4.2.1 Diverse Decision Strategies Across Visualizations
Strategies in Both Visualizations: People First vs. Distance First. Participants primarily adopted one of two strategies: (1) prioritizing fewer people in line, then considering walking distance; or (2) prioritizing shorter distance, then checking queue length. In the Minimap conditions, participants were almost evenly split between these two strategies, with roughly half employing a “people first” (15 out of 32) and the other half a “distance first” (14 out of 32). By contrast, in the AR conditions, a slight preference emerged for distance-first strategies (15 out of 32), while fewer participants initially considered queue lengths (11 out of 32). The remaining participants described balancing both factors equally without a dominant criterion.
AR: Unique Heuristics and Motivations. Whereas strategy selection in the Minimap conditions appeared to depend largely on participants’ spatial skills, the AR conditions introduced distinct perceptual heuristics influencing participants to prioritize distance:
  • Egocentric Spatial Imagery. AR see-through’s egocentric view led participants to vividly imagine physical navigation, especially complex pathways with hallway turns or floor changes. For example, P9 noted, “It’s easier to draw or imagine a line from where I am to the coffee machine…” This detailed mental imagination caused them to weigh distance more heavily in their decisions.
  • Visual Proximity Bias. Participants described a misleading visual proximity effect, where certain targets appeared visually closer (Figure 7a) due to the see-through nature of the AR visualizations, even if actual walking paths were longer (Figure 7b). As P8 noted, “I think this way is shorter but it’s actually longer.” This illusion occasionally influenced participants to overweight distance.
  • Occlusion Challenges. Participants encountered difficulties estimating queue lengths due to occlusion in the AR visualizations, prompting reliance on distance estimates instead of people counts.
Summary. Compared to the Minimap conditions, participants in the AR conditions showed a stronger preference for spatial reasoning (distance) in their decision-making. This suggests that the embedded AR-see through visualization inherently emphasizes spatial perception, which, however, can be a double-edged sword that introduces perceptual biases and occlusion challenges.
4.2.2 Increased Trust in AI Suggestions with AR See-Through
The visualizations impacted participants’ trusted and relied on AI.
Minimap: Make Own Choice First, Then Consult AI. In the Minimap conditions, most participants (22 out of 32) preferred first to form their own judgment and then cross-check against the AI suggestion (e.g., “I use it just to double-check my decision”, P22). Participants reported feeling more confident in their own judgment when using the Minimap visualization because it presented the information required to make informed choices. They typically consulted the AI in cases of uncertainty or indecision: “If I felt stuck between two locations, I would check the AI to break the tie” (P20).
AR: AI as a Starting Anchor. Participants expressed greater trust and reliance on AI suggestions in the AR conditions, frequently using the AI suggestion as an initial reference or anchor for their decisions. Several participants (e.g., P6, P11) mentioned that AI suggestions helped narrow the decision space and enabled quicker comparisons. Participants provided several reasons for trusting AI more with AR:
  • Perceptual and navigation challenges. Participants found distance estimation (e.g., P2), queue length assessment (e.g., P4, P5), and locating navigation paths more effortful in AR (e.g., P12). Thus, AI suggestions offered a useful cognitive shortcut, aligning with findings from prior research on cognitive load and decision-making [3].
  • Unfamiliarity with AR See-Through visualizations. Some participants mentioned that unfamiliar with the AR See-Through visualizations increased their reliance on AI guidance (e.g. P31).
  • Enhanced Realism and Embodiment of AI suggestions. Interestingly, several participants indicated that the realistic, embodied way AI suggestions appeared in the AR See-Through increased their sense of trust and acceptance (e.g., P20, P21).
Summary. compared to the Minimap, the AR visualization led to greater participant trust in and reliance on AI suggestions. Participants attributed this shift to greater perceptual complexity, unfamiliarity with the AR visualizations, and the more realistic presentation of AI.
5 Discussion
In this section, we synthesize the key lessons learned, design implications, and promising future directions to advance human-AI collaboration in spatial tasks in AR.
5.1 Perceptual Challenges of AR-Embedded Visualizations Can Lead to AI Over-Reliance
Our findings suggest that perceptual challenges of embedded visualizations in AR – such as occlusion, large canvas space, and spatial estimation difficulty – can increase users’ reliance on AI suggestions, even when those suggestions are suboptimal. This over-reliance does not appear to stem from blind trust in AI, but rather from the inherent complexity of perceiving and evaluating spatial options in large-scale, dynamic physical environments. While prior research has discussed AR’s perceptual limitations [18], our study directly connects these challenges to AI reliance patterns in time-sensitive spatial decision-making. This highlights the need for better embedded visualization designs that support spatial reasoning, especially in large, dynamic environments.
Design Implications. Providing AI support in AR environments is not inherently beneficial. It can be neutral or even detrimental when perceptual challenges make it harder to verify AI suggestions. These findings underscore the need to rethink how AI assistance is integrated into AR, rather than directly porting AI designs from desktop or 2D interfaces. Designers should explicitly consider how spatial perception difficulties affect users’ ability to evaluate AI output.
Future Studies. In our study, situational data (e.g., queue length) was visualized using human-shaped glyphs. A natural next step is to compare visual encodings (e.g., glyphs vs. numeric labels) to evaluate how data representation influences AI reliance. Additionally, examining how different levels of spatial complexity (e.g., number of targets, layout density) modulates trust and reliance in AR-embedded visualizations would offer deeper insight into AR-AI interaction design.
5.2 Cognitive Biases Triggered by AR-Embedded Visualizations Can Promote Inappropriate AI Reliance
Beyond perceptual difficulty, our qualitative findings point to cognitive biases that are unique to AR-embedded visualizations. Participants reported visual proximity illusions and increased trust in AI suggestions due to their spatial embodiment in the environment. These effects align with prior research on spatial cognition biases [60] and trust in embodied agents [37].
Although cognitive biases have been studied in traditional visualization contexts [15, 82], their implications in embodied, spatial, and AI-augmented AR systems remain underexplored. As Padilla et al. [54] suggested, when systems are designed with such biases in mind, users can benefit from Type 1 (intuitive) thinking—particularly under time pressure. Our findings highlight a fertile area for interdisciplinary research at the intersection of spatial cognition, embedded visualizations, and human-AI decision-making.
Design Implications. Embedded visualizations in AR should explicitly account for and possibly counteract spatial cognitive biases. For example, visual proximity illusions may be mitigated through alternative depth cues, or warnings when direct-line views differ from walking paths. Designers might also consider using less embodied, realistic representation for AI suggestions when the risk of over-reliance is high.
Future Studies. While spatial AR studies are often limited by physical constraints of real-world spaces, the increasing accessibility of AR and AI-integrated devices makes spatial cognition experiments increasingly feasible. Future research should investigate how AR-induced biases interact with varying levels of AI accuracy or explanation quality, particularly in high-stakes or time-sensitive scenarios. This represents a promising interdisciplinary frontier spanning visual computing, cognitive science, and human-computer interaction.
5.3 Aligning the Strengths of AR-Embedded Visualizations with Spatial Tasks
In our study, while the AR see-through did not outperform the 2D minimap on decision accuracy, it significantly reduced post-trial pointing errors, suggesting improved spatial mapping and memory. This distinction suggests that AR-embedded visualizations may be more beneficial in continuous, action-based spatial workflows rather than in isolated decision-making tasks.
In high-stakes domains like emergency response, selecting a correct location is only the first step – what follows is action. Misinterpreting a target’s location after selection can be consequential. For instance, a user might correctly choose an exit from a wall-mounted map but walk the wrong way due to reference-frame confusion. AR-embedded visualizations mitigates this risk by spatially anchoring information in the environment. Thus, rather than emphasizing the weakness of AR-embedded visualizations, our findings illustrate its potential strengths in embodied spatial tasks and highlight when and where AR-embedded visualizations are best applied.
Design Implications. AR-embedded visualizations may be most effective when tightly coupled with tasks involving physical action, rather than static decision-making. Designers should consider the viewer’s movement and real-time orientation when designing the embedded visualizations. Emerging work on “Visualization in Motion” [94] and “First-Person View Visualization” [34] provides valuable guidance in this direction. In addition, AR systems should support smooth transitions between embedded visualizations and traditional 2D visualizations in the headset depending on task context and user needs.
Future Studies. In our study, participants selected targets verbally. A natural follow-up would involve embodied selection, e.g., walking to or pointing at targets, to examine whether the spatial support of AR-embedded visualizations becomes more critical under physical interaction. We hypothesize that in such embodied tasks, participants using 2D maps will incur greater cognitive load in spatial mapping, possibly diminishing their ability to evaluate targets effectively. Future work should explore how task modality (e.g., static vs. embodied) interacts with visualization and AI support in complex environments.
5.4 AR-Embedded Visualizations May Lower Cognitive Engagement Compared with 2D Minimap
In our study, there was no significant difference in response time between Minimap+NoAI and AR+NoAI once initial search time was excluded. However, participants in AR+AI made decisions more quickly yet achieved lower overall accuracy than those in Minimap+AI. This pattern aligns with reliance on heuristic judgments – often described as Type 1 thinking [12, 22]. Together with the high self-reported confidence, these results suggest a lack of deeper cognitive engagement: once AI assistance was available, participants quickly felt confident enough in their intuitive judgments, rather than engaging in more analytical (Type 2) thinking. These findings imply that AR-embedded visualizations, by offering a highly immersive or embodied experience, may inadvertently reduce users’ motivation to scrutinize information.
Design Implications. It remains unclear whether this phenomenon is specific to our task design, a broader characteristic of AR see-through interfaces, or a more general cognitive shift from 2D to AR-embedded. Design strategies that prompt more deliberation – such as “cognitive forcing functions” [10] – could be incorporated to encourage deeper analytic thinking in AR environments.
Future Studies. Further research is needed to determine whether AR-embedded visualizations consistently reduces cognitive motivation compared to 2D visualizations, and to explore methods for mitigating AR-induced cognitive shortcuts across diverse tasks, visualization designs, and populations.
5.5 Study Limitations
Our study offers an initial examination of AI-aided spatial decision-making with AR see-through visualizations. Although guided by prior research on AR visual systems for spatial tasks [91, 90], we acknowledge that our system design does not represent a fully optimized or universal solution. To ensure experimental control, we adopted fixed parameters, such as a discrete set of spatial arrangements, an AI accuracy of 75%, and a 20-second decision window, that may not capture the full complexity of high-stakes, real-world environments or more advanced AI systems. Consequently, our findings are most transferable to tasks resembling the indoor, time-sensitive context described in Section 2.1.
Moreover, our participant pool primarily comprised well-educated individuals, which may not reflect the behavior of specialized user groups (e.g., emergency responders or security professionals). Finally, as is common in spatial cognition research [75], our sample size was modest; although sufficient for controlled experimentation, a larger and more diverse sample would strengthen external validity. We envision this work as a stepping stone for future studies to extend and refine AR-based decision support across broader tasks and populations.
6 Conclusion
In this paper, we investigated the impact of AR see-through visualizations on human-AI reliance in spatial decision-making under time constraints by comparing them with traditional 2D minimaps. Our user study (N = 32) revealed that AR see-through displays tend to increase over-reliance on AI during spatial decision-making. This effect is primarily attributable to perceptual challenges, such as occlusion and spatial estimation difficulties, as well as AR-specific biases like visual proximity illusions and the influence of high-fidelity visual representations. Nonetheless, AR exhibits unique strengths by enhancing spatial reasoning, as evidenced by significantly reduced pointing errors that suggest improved spatial mapping and navigation. We hope that our findings and design implications will guide future research and the development of solutions that better leverage AR’s advantages to improve human-AI collaboration in high-stakes, spatial contexts.